Please helpme
A maternity hospital located at the heart of the city is providing nursery kit items for expectant motjer clients. The common color of these item either pink (for female babies) or blue (fpr male babies). Thr hospital management concern is the number of each colored item they have to prepare to meet the demand of their clients. On the basis of their previous years experience, the management estimates that the probability thay any one mother will have female baby and will be using pink items is 60. Determine the probability distribution from 20 mothers with n=20 and p=60 with the table below. The table is numbers 4 - 11

Answer by rothauserc(4718)

You can use the binomial probability formula,
:
Probability (P) (k successes in n trials) = nCk * p^k ^ (1-p)^(n-k), where nCk = n!/(k! * (n-k)!)
:
For this problem, n = 20, p = 0.60
:
For example, k = 4
:
P (4 pink nursery kits needed out of 20) = 20C4 * (0.60)^4 * (1-0.60)^(20-4) = 0.0003
:
You can use your calculator for the rest
: